The condensed income statement for the Carla Vista and Paul partnership for 2020 is as follows.

Carla Vista and Paul Company
Income Statement
For the Year Ended December 31, 2020

Sales (270,000 units) $1,350,000
Cost of goods sold 900,000
Gross profit 450,000
Operating expenses
Selling $270,000
Administrative 189,000
459,000
Net loss $(9,000 )


A cost behavior analysis indicates that 75% of the cost of goods sold are variable, 42% of the selling expenses are variable, and 40% of the administrative expenses are variable.Compute the break-even point in total sales dollars for 2020. (Round intermediate calculations to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 2,520.)

Break-even point in dollars $
Answer #1

Variable cost = 900000*75% + 270000*42% + 189000*40% = 864000

Fixed cost = 900000+459000-864000 = 495000

Contribution margin = (1350000-864000) /1350000 = 36%

Break even point in dollars = fixed cost / contribution margin

= 495000 / 36% = 1375000

Break even point in units = 270000*1375000 /1350000= 275000 units